OpenStudy (anonymous):

9x(5x^2) is this 45x^2 orr 45x^3? I'm a bit confused :(

OpenStudy (diyadiya):

\[\large x^m \times x^n = x^{m+n}\] So \[x^{1} \times x^2 = ?\]

OpenStudy (anonymous):

1 * 2 = 2 So idk if i multiply them or if i add them 1 + 2 = 3

OpenStudy (diyadiya):

\[x^1 + x^2 = x^{1+2} \] We add the powers (:

OpenStudy (anonymous):

(((: Thank you soo much does that go for both sides example if it had of been 9x^2(5x) would we have gotten the same answer?

OpenStudy (diyadiya):

Yes Same answer \[x^2 \times x^2 =x^{2+1}=x^3\] and 9 x 5 =45 5 x 9 =45 :)

OpenStudy (anonymous):

Thanks!! ((((: